​For a first order (A)-------products , the concentration of A changes from 0.1 M to 0.025 M in 40 minutes. the rate of reaction when the concentration of A is 0.01 M is.....??
plz. provide detailed solution
 

Grade:12

1 Answers

Sakshi
8 years ago
For the first order reaction, one can write,
A=A0e-kt
thus, one can write, 0.025=0.1 * e-k*40
on solving, k= 0.03465

now, also one can write, Rate= k[A]= 0.035*0.01=0.00035min-1